OpenStudy (anonymous):

Identify the point (x1, y1) from the equation: y – 7 = 3(x + 2) (–7, 2) (7, –2) (–2,7) (2, –7) I know the answer is 2,7 but which one is negative???? that's what confuses me

OpenStudy (anonymous):

Its easy my dear friend.... just substitute every point in the given eqution... if any point satisfies the eqution then that point is (x1,y1)

OpenStudy (nurali):

i think (–2,7) 7-7 = 3(-2+2) 0=0

OpenStudy (anonymous):

So the point slope equation is \[y-y _{1}=m(x-x _{1})\] You can read off the solution : y-7 on the left hand side, thus y1=7 3(x+2) on the right hand side, thus x1=-2 (I hope it is clear why -(-)=+)

OpenStudy (anonymous):

Another way to identify is.... if we place 7 in y left handside part will become zero then -2 in place of x then right hand side part will become zero so, answer must be (-2,7)